OpenStudy (anonymous):

Change (3, 2π) into rectangular form.

OpenStudy (anonymous):

\[x=3\cos 2\pi,y=3\sin 2\pi~ squaring ~ and~ adding,we~ get ~?\]

OpenStudy (anonymous):

well I got (3,0) as my coordinates, but idk if thats right

OpenStudy (anonymous):

\[x^2+y^2=3^2=9\]

OpenStudy (anonymous):

ok, what's next?

OpenStudy (anonymous):

(9, 1) ???

OpenStudy (anonymous):

i have completed the solution.

OpenStudy (anonymous):

\[x^2=3^2\cos ^22\pi \] similarly for y^2 and then add

OpenStudy (anonymous):

wait you mean, x^2+y^2=3^2=9 is the answer?

OpenStudy (anonymous):

wait I don't get it

OpenStudy (anonymous):

like for this problem Change (2, 20°) into rectangular form. I got (1.9, 0.68)

OpenStudy (anonymous):

was your answer correct?

OpenStudy (anonymous):

i don't know since I haven't submitted it

OpenStudy (anonymous):

how you did it ?

OpenStudy (anonymous):

you are on complex numbers or changing polar to rectangular co-ordinates.

OpenStudy (anonymous):

x=3cos 2 pi=3*-1=-3 y=3 sin 2 pi=3*0=0 rectangular coordinates are (3,0) if you want equation in rectangular form ,that i have already given.

OpenStudy (anonymous):

corection x=3 cos 2 pi=3*1=3

OpenStudy (anonymous):

x=2 cos 20=1.9 y=2 sin 20=0.68